# UniTable: Towards a Unified Framework for Table Structure Recognition via Self-Supervised Pretraining

### v1

@@ Abstract
Tables convey factual and quantitative data with implicit conventions created by humans that are often challenging for machines to parse. Prior work on table structure recognition (TSR) has mainly centered around complex task-specific combinations of available inputs and tools. We present UniTable, a training framework that unifies both the training paradigm and training objective of TSR. Its training paradigm combines the simplicity of purely pixel-level inputs with the effectiveness and scalability empowered by self-supervised pretraining (SSP) from diverse unannotated tabular images. Our framework unifies the training objectives of all three TSR tasks — extracting table structure, cell content, and cell bounding box (bbox) — into a unified task-agnostic training objective: language modeling. Extensive quantitative and qualitative analyses highlight UniTable’s state-of-the-art (SOTA) performance on four of the largest TSR datasets To promote reproducible research, enhance transparency, and SOTA innovations, we open-source our code at https://github.com/poloclub/unitable and release the first-of-its-kind Jupyter Notebook of the whole inference pipeline, fine-tuned across multiple TSR datasets, supporting all three TSR tasks.
@@ 1 Introduction
Tables are ubiquitous in documents, as they serve to summarize factual and quantitative data — information that is cumbersome to describe in text but nevertheless crucial (Göbel et al., 2013; Gao et al., 2019). Due to the implicit conventions used by humans in creating tables, the representations within tables are often challenging for machines to parse. Even the milestone vision-language models (VLM), GPT-4V(vision) (Yang et al., 2023), still struggles with various document-related tasks including TSR (Cui et al., 2023). GPT-4V tends to omit content in large tables and performs worse when faced with complex tables, e.g., spanning or empty cells and uneven text distributions (Shi et al., 2023).
@@
Prior work on TSR has mainly centered around complex task-specific combinations of available inputs and tools. Typically, table structure was predicted by an image-to-text pipeline (Zhong et al., 2020) and cell bbox was predicted by a detection head, e.g., Faster R-CNN (Smock et al., 2022) or DETR (Nassar et al., 2022). The assumption regarding predicting cell content varies: some studies assume the presence of a portable document format (PDF) accompanying the tabular image (Nassar et al., 2022), while others rely on external text line detection and text recognition models (Ye et al., 2021; Huang et al., 2023). However, training generic Transformers under a single language modeling objective, i.e., predicting the next token, for diverse tasks has achieved remarkable success across diverse tasks in language (Radford et al., 2019; Brown et al., 2020), vision (Chen et al., 2021, 2022), and vision-and-language domains (Achiam et al., 2023). We wonder whether understanding visual language in tables, i.e., extracting table structure, cell content, and cell bbox from tabular images, can be seamlessly integrated into the language modeling training framework. This integration is challenging because: (1) the fusion of vision and language in tabular images demands high-fidelity reading and rich high-level representations (Lee et al., 2023); (2) the large amount of unannotated tabular images in practice cannot be leveraged by existing supervised learning approaches (Nassar et al., 2022; Ma et al., 2023); (3) the diverse output of table-related tasks are typically addressed by task-specific models (Huang et al., 2023); and (4) direct application of a linear projection Transformer results in the significant performance drop (Peng et al., 2023), leading prior work to exclusively employ a hybrid convolutional neural network (CNN)-Transformer architecture. We resolve the above challenges by proposing UniTable, a training framework that unifies both training paradigm and training objective of TSR and make the following major contributions (Fig. 1):
@@
Figure 1: UniTable, a training framework that unifies both training paradigm and training objective of TSR. In UniTable, the visual encoder is self-supervised pretrained and then finetuned along with the task decoder on supervised datasets. UniTable unifies the training objectives of all three TSR tasks — extracting table structure, cell bbox, and cell content — into a unified task-agnostic training objective: language modeling. With UniTable, the user inputs a tabular image and obtains the corresponding digitalized table in HTML.
@@
    UniTable’s training paradigm combines the simplicity of purely pixel-level inputs with the effectiveness and scalability empowered by SSP from diverse unannotated tabular images. Specifically, UniTable unifies the training paradigm for TSR: pretraining the visual encoder by predicting the masked tabular images in a self-supervised manner and finetuning the visual encoder along with the task decoder on supervised datasets. With UniTable, the table structure prediction on SynthTabNet (Nassar et al., 2022), a comprehensive dataset with 600k tables across finance, marketing, and academia in both dense and sparse format, achieves the SOTA 99.18% when self-supervised pretrained on 2M images, significantly lifting the original accuracy of 84.04% when trained from scratch using a linear projection Transformer. Owing to the powerful SSP, UniTable has also successfully mitigated the performance drop caused by replacing the CNN backbone with the linear projection.
@@
    UniTable unifies the training objectives of all three TSR tasks — extracting table structure, cell content, and cell bbox — into a unified task-agnostic training objective: language modeling. Specifically, the input to our model is an image in the form of raw pixels only and the output is text in the form of token sequences, and the training objective is language modeling. UniTable’s SOTA performance on the FinTabNet dataset (Zheng et al., 2021) demonstrates our approach’s generalizability to the PDF input modality as we can simply convert PDF to images. Our framework also enables us to leverage the power of SSP on large-scale unannotated tabular images as all models are finetuned from SSP. UniTable’s unified training objective applies to both linear projection Transformer and hybrid CNN-Transformer architectures conventionally used in TSR.
@@
    Extensive quantitative and qualitative analyses highlight UniTable’s SOTA performance on four of the largest TSR datasets: ICDAR 2019 B2 Modern (Gao et al., 2019), PubTabNet (Zhong et al., 2020), FinTabNet (Zheng et al., 2021), and SynthTabNet (Nassar et al., 2022). Our visualization of the visual tokens reveals the key reason for why SSP works — the visual semantics captured by the visual codebook show a fine-grained categorization to represent the implicit human conventions used when creating the tables.
@@
    Open-source code and UniTable in practice. To promote reproducible research, enhance transparency, SOTA innovations, and facilitate fair comparisons in our domain as tables are a promising modality for representation learning, we open-source our code at https://github.com/poloclub/unitable. We provide all the details regarding training, validation, testing, and ablation studies. To enable users to easily try UniTable on their own tabular images and obtain fully digitized HTML tables, we release the first-of-its-kind Jupyter Notebook of the whole inference pipeline, fine-tuned across multiple TSR datasets, supporting all three TSR tasks.
@@ 2.1 Task Definition
The goal of TSR is to translate the input tabular image $\mathbf{I}$ into a machine-readable sequence $\mathbf{T}$, which typically consists of table structure $\mathbf{S}$, table cell bbox $\mathbf{B}$, and table cell content $\mathbf{C}$. Structure $\mathbf{S}=[s_{1},...,s_{m}]$ is a sequence of tokenized HTML table tags $s$, cell bbox $\mathbf{B}=[\mathbf{b}_{1},...,\mathbf{b}_{n}]$ is a sequence of bboxes defined by $\mathbf{b}=(x_{min},y_{min},x_{max},y_{max})$, and cell content $\mathbf{C}=[c_{1},...,c_{n}]$ comprises the content within each cell in reading order. Note that the sequence length of $\mathbf{B}$ and $\mathbf{C}$ are the same, but shorter than $\mathbf{S}$ as the HTML tags contain both empty and non-empty cells. Since each cell $i$ is defined by a single bbox, $c_{i}$ can have either single or multiple text lines.
@@ 2.2 Model Architecture
tsr model architecture has two modules: visual encoder and task decoder. The visual encoder extracts features from input $\mathbf{I}$, and the task decoder predicts $\mathbf{T}$. For the visual encoder, prior work employed either an off-the-shelf CNN backbone, e.g., ResNets and ResNet variants (He et al., 2016), or hybrid CNN-Transformer architecture. EDD (Zhong et al., 2020) explored five different ResNet-18 variants, TableMaster (Lu et al., 2021) combined residual blocks with multi-aspect global context attention, TableFormer (Nassar et al., 2022) connected a ResNet-18 with Transformer encoder layers, and VAST (Huang et al., 2023) adopted the first four stages of a ResNet-31. These convolution layers cannot be replaced because a direct employment of a vanilla Transformer with linear projection leads to a significant performance drop (Peng et al., 2023). However, linear projection, that divides image into patches, is a widely used input image processor in SOTA vision Transformers (Dosovitskiy et al., 2020; Liu et al., 2021b), VLMs (Li et al., 2022; Liu et al., 2023) and multi-modal models (Mizrahi et al., 2023). Recent work on visual language understanding has also successfully adopted the linear projection (Kim et al., 2022; Lee et al., 2023). Thus, to avoid having a separate architecture design solely for the table domain, we aim to keep the linear projection Transformer, and mitigate the performance gap by SSP. Sec. 3 shows that pretraining the visual encoder in a self-supervised manner significantly helps the model learn how to parse the table structure and achieves performance even higher than architectures with convolutions.
3 A Unified Framework for Pretraining and Finetuning TSR Models
We pretrain the visual encoder by predicting the masked tabular images in a self-supervised manner and finetune the visual encoder along with the task decoder using the supervised dataset for each task, as shown in Fig. 1. Sec. 3.1 introduces the SSP of the visual encoder, and Sec. 3.2 details how we unify all TSR tasks and finetune with the pretrained visual encoder.
@@ 3.1 Self-Supervised Pretraining of the Visual Encoder
Image tokens. Define a visual codebook in the latent space $Z\in\mathbb{R}^{K\times D}$, representing $K$ entries of visual tokens $z_{i}\in\mathbb{R}^{D}$. The visual codebook is trained in such a way that an input image, once it is embedded into an image grid, each pixel on the embedded image grid can be substituted with a visual token from the codebook. Decoding this modified image grid will then reconstruct the input image. We use Vector Quantized-Variational AutoEncoder (VQ-VAE(Van Den Oord et al., 2017) to train the visual codebook. Specifically, the tabular image $I$ is tokenized into discrete tokens $z$ after passing through the encoder $q_{\phi}(z|I)$, and the decoder $p_{\psi}(I|z)$ takes these discrete tokens and rebuilds the original image. The training objective of VQ-VAE is to maximize the $\mathbb{E}_{z\sim q_{\phi}(z|I)}[\log p_{\psi}(I|z)]$ with respect to $\phi$ and $\psi$. The training is non-differentialble due to the categorical distribution in selecting visual tokens. Thus, we use Gumbel-Softmax (Jang et al., 2016) as a reparameterization trick following DALL-E (Ramesh et al., 2021). We have trained the VQ-VAE on 1M and 2M tabular images, where $K=8192$ for 1M, and $K=16384$ for 2M.
Image patches. Given an input tabular image $I\in\mathbb{R}^{H\times W\times C}$, the linear projection divides $I$ into a sequence of flattened 2D patches $I_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $C$ is the number of channels, $(P,P)$ is the size of each image patch, and $N=HW/P^{2}$ is the number of patches. It is implemented by a kernel $P\times P$, stride $P$ convolution. We set $P=16$ and $I\in\mathbb{R}^{448\times 448\times 3}$, thus the sequence length of the image patches is $28\times 28$. Approximately 40% of the sequence is replaced with a masked token and the pretraining objective is to maximize the log-likelihood of the visual tokens of the masked region given the unmasked region. The groundtruth visual tokens are provided by the codebook of image tokenizer.
The pretraining task is inspired by the success of masked language (Devlin et al., 2018) and natural image (Bao et al., 2021) modeling, but our work differs in the following ways:
@@ 3.2 Unified Finetuning Framework
Prior work employed a task-specific decoder for each task, where $\mathbf{S}$ was predicted by an image-to-text pipeline and $\mathbf{B}$ was predicted by a detection head, e.g., Faster R-CNN (Smock et al., 2022) or DETR (Nassar et al., 2022). The assumption on predicting $\mathbf{C}$ varies: some assume a PDF is always accompanying the tabular image (Nassar et al., 2022), others rely on external text line detection and text recognition models (Ye et al., 2021; Huang et al., 2023). We aim to provide a unified task-agnostic training framework, where the input to our model is an image in the form of raw pixels only and the output is text in the form of token sequences. This setting is also generalizable to PDF input modality as we can simply take a screenshot of the PDF. The framework also enables us to leverage the visual encoder pretrained on the unannotated tabular images.
@@
Table cell bbox $\mathbf{B}$. Each cell bbox $\mathbf{b}$ are four continuous coordinates, which are not naturally expressed as discrete tokens. Inspired by Pix2Seq (Chen et al., 2021), we discretize the coordinate into an integer between 0 and image size. The two directions of an image share the same vocabulary. Since we need to predict all bboxes within a tabular image, each quantized bbox is concatenated together in reading order: from left to right and top to bottom. This formulation completes the quantization and serialization of all bboxes into a sequence of discrete tokens. At inference time, we de-serialize the predicted sequence into groups of four tokens.
Table cell content $\mathbf{C}$. After predicting all bbox coordinates, we only need to perform optical character recognition (OCR) on the image region within each bbox. Note that each cell is defined by a single bbox, thus the cell content can have single or multiple lines of text. In the training stage, the model is trained on a mixture of single line and multi-line dataset. At inference time, we parse all text simultaneously as each cell bbox is independent. Finally, we insert the cell content back into the non-empty cells <td>[]</td> or >[]</td> as the reading order is already preserved in both S and B. We use WordPiece tokenizer (Wu et al., 2016) with charater-level granularity since OCR requires the model to read instead of understanding the semantics, which significantly reduces the total vocabulary size to less than 6k.
@@
Table 1: UniTable outperforms prior methods and achieves SOTA on four out of five largest publicly available TSR datasets across all available tasks. Our method is trained with a task-agnostic language modeling loss and does not rely on external PDF for text extraction and bbox post-processing.

### v2

# UniTable: Towards a Unified Framework for Table Recognition via Self-Supervised Pretraining

@@ Abstract
Tables convey factual and quantitative data with implicit conventions created by humans that are often challenging for machines to parse. Prior work on table recognition (TR) has mainly centered around complex task-specific combinations of available inputs and tools. We present UniTable, a training framework that unifies both the training paradigm and training objective of TR. Its training paradigm combines the simplicity of purely pixel-level inputs with the effectiveness and scalability empowered by self-supervised pretraining (SSP) from diverse unannotated tabular images. Our framework unifies the training objectives of all three TR tasks — extracting table structure, cell content, and cell bounding box (bbox) — into a unified task-agnostic training objective: language modeling. Extensive quantitative and qualitative analyses highlight UniTable’s state-of-the-art (SOTA) performance on four of the largest TR datasets. UniTable’s table parsing capability has surpassed both existing TR methods and general large vision-language modelss, e.g., GPT-4o, GPT-4-turbo with vision, and LLaVA. Our code is publicly available at https://github.com/poloclub/unitable, featuring a Jupyter Notebook that includes the complete inference pipeline, fine-tuned across multiple TR datasets, supporting all three TR tasks.
@@ 1 Introduction
Tables are ubiquitous in documents, as they serve to summarize factual and quantitative data — information that is cumbersome to describe in text but nevertheless crucial [10, 8]. Due to the implicit conventions used by humans in creating tables, the representations within tables are often challenging for machines to parse. Even the milestone VLM, GPT-4o [29], GPT-4V(vision) [39] and LLaVA [21], still struggles with various document-related tasks including TR. GPT-4V tends to omit content in large tables and performs worse when faced with complex tables, e.g., spanning or empty cells and uneven text distributions [35].
@@
Prior work on TR has mainly centered around complex task-specific combinations of available inputs and tools. Typically, table structure was predicted by an image-to-text pipeline [42] and cell bbox was predicted by a detection head, e.g., Faster R-CNN [36] or DETR [28]. The assumption regarding predicting cell content varies: some studies assume the presence of a portable document format (PDF) accompanying the tabular image [28], while others rely on external text line detection and text recognition models [40, 12]. However, training generic Transformers under a single language modeling objective, i.e., predicting the next token, for diverse tasks has achieved remarkable success across diverse tasks in language [32, 3], vision [4, 5], and vision-and-language domains [1]. We wonder whether understanding visual language in tables, i.e., extracting table structure, cell content, and cell bbox from tabular images, can be seamlessly integrated into the language modeling training framework. This integration is challenging because: (1) the fusion of vision and language in tabular images demands high-fidelity reading and rich high-level representations [15]; (2) the large amount of unannotated tabular images in practice cannot be leveraged by existing supervised learning approaches [28, 26]; (3) the diverse output of table-related tasks are typically addressed by task-specific models [12]; and (4) direct application of a linear projection Transformer results in the significant performance drop [30], leading prior work to exclusively employ a hybrid convolutional neural network (CNN)-Transformer architecture. We resolve the above challenges by proposing UniTable, a training framework that unifies both training paradigm and training objective of TR and make the following major contributions (Fig. 1):
@@
Figure 1: UniTable, a training framework that unifies both training paradigm and training objective of TR. In UniTable, the visual encoder is self-supervised pretrained and then finetuned along with the task decoder on supervised datasets. UniTable unifies the training objectives of all three TR tasks — extracting table structure, cell bbox, and cell content — into a unified task-agnostic training objective: language modeling. With UniTable, the user inputs a tabular image and obtains the corresponding digitalized table in HTML.
Figure 2: The table parsing capability of UniTable has surpassed that of general large VLMs, e.g., GPT-4o, GPT-4-turbo with vision, and LLaVA. For complex tables that include multiple spanning cells, UniTable can successfully reconstruct the table in HTML, whereas general large VLMs fail in various aspects.
@@
    UniTable’s training paradigm combines the simplicity of purely pixel-level inputs with the effectiveness and scalability empowered by SSP from diverse unannotated tabular images. Specifically, UniTable unifies the training paradigm for TR: pretraining the visual encoder by predicting the masked tabular images in a self-supervised manner and finetuning the visual encoder along with the task decoder on supervised datasets. With UniTable, the table structure prediction on SynthTabNet [28], a comprehensive dataset with 600k tables across finance, marketing, and academia in both dense and sparse format, achieves the SOTA 99.18% when self-supervised pretrained on 2M images, significantly lifting the original accuracy of 84.04% when trained from scratch using a linear projection Transformer. Owing to the powerful SSP, UniTable has also successfully mitigated the performance drop caused by replacing the CNN backbone with the linear projection.
@@
    UniTable unifies the training objectives of all three TR tasks — extracting table structure, cell content, and cell bbox — into a unified task-agnostic training objective: language modeling. Specifically, the input to our model is an image in the form of raw pixels only and the output is text in the form of token sequences, and the training objective is language modeling. UniTable’s SOTA performance on the FinTabNet dataset [41] demonstrates our approach’s generalizability to the PDF input modality as we can simply convert PDF to images. Our framework also enables us to leverage the power of SSP on large-scale unannotated tabular images as all models are finetuned from SSP. UniTable’s unified training objective applies to both linear projection Transformer and hybrid CNN-Transformer architectures conventionally used in TSR.
@@
    Extensive quantitative and qualitative analyses highlight UniTable’s SOTA performance on four of the largest TR datasets: ICDAR 2019 B2 Modern [8], PubTabNet [42], FinTabNet [41], and SynthTabNet [28]. UniTable’s table parsing capability has surpassed both existing TR methods and general large VLMs (Fig. 2), e.g., GPT-4o, GPT-4V(ision), and LLaVA. Due to our unified language modeling framework formulation, we discover three types of previously unacknowledged inconsistencies in the groundtruth annotations of PubTables-1M  [36], one of the largest TSR datasets, accounting for more than 53.10% of its training set. Our visualization of the visual tokens reveals the key reason for why SSP works — the visual semantics captured by the visual codebook show a fine-grained categorization to represent the implicit human conventions used when creating the tables.
@@
    Open-source code and UniTable in practice. To promote reproducible research, enhance transparency, SOTA innovations, and facilitate fair comparisons in our domain as tables are a promising modality for representation learning, we open-source our code (anonymized) [^1] https://github.com/poloclub/unitable . We provide all the details regarding training, validation, testing, and ablation studies. To enable users to easily try UniTable on their own tabular images and obtain fully digitized HTML tables, we release the first-of-its-kind Jupyter Notebook of the whole inference pipeline, fine-tuned across multiple TR datasets, supporting all three TR tasks.
@@ 2.1 Task Definition
The goal of TR is to translate the input tabular image $\mathbf{I}$ into a machine-readable sequence $\mathbf{T}$, which typically consists of table structure $\mathbf{S}$, table cell bbox $\mathbf{B}$, and table cell content $\mathbf{C}$. Structure $\mathbf{S}=[s_{1},...,s_{m}]$ is a sequence of tokenized HTML table tags $s$, cell bbox $\mathbf{B}=[\mathbf{b}_{1},...,\mathbf{b}_{n}]$ is a sequence of bboxes defined by $\mathbf{b}=(x_{min},y_{min},x_{max},y_{max})$, and cell content $\mathbf{C}=[c_{1},...,c_{n}]$ comprises the content within each cell in reading order. Note that the sequence length of $\mathbf{B}$ and $\mathbf{C}$ are the same, but shorter than $\mathbf{S}$ as the HTML tags contain both empty and non-empty cells. Since each cell $i$ is defined by a single bbox, $c_{i}$ can have either single or multiple text lines.
@@ 2.2 Model Architecture
tsr model architecture has two modules: visual encoder and task decoder. The visual encoder extracts features from input $\mathbf{I}$, and the task decoder predicts $\mathbf{T}$. For the visual encoder, prior work employed either an off-the-shelf CNN backbone, e.g., ResNets and ResNet variants [11], or hybrid CNN-Transformer architecture. EDD [42] explored five different ResNet-18 variants, TableMaster [24] combined residual blocks with multi-aspect global context attention, TableFormer [28] connected a ResNet-18 with Transformer encoder layers, and VAST [12] adopted the first four stages of a ResNet-31. These convolution layers cannot be replaced because a direct employment of a vanilla Transformer with linear projection leads to a significant performance drop [30]. However, linear projection, that divides image into patches, is a widely used input image processor in SOTA vision Transformers [7, 22], VLMs [16, 20] and multi-modal models [27]. Recent work on visual language understanding has also successfully adopted the linear projection [14, 15]. Thus, to avoid having a separate architecture design solely for the table domain, we aim to keep the linear projection Transformer, and mitigate the performance gap by SSP. Sec. 3 shows that pretraining the visual encoder in a self-supervised manner significantly helps the model learn how to parse the table structure and achieves performance even higher than architectures with convolutions.
3 A Unified Framework for Pretraining and Finetuning TR Models
We pretrain the visual encoder by predicting the masked tabular images in a self-supervised manner and finetune the visual encoder along with the task decoder using the supervised dataset for each task, as shown in Fig. 1. Sec. 3.1 introduces the SSP of the visual encoder, and Sec. 3.2 details how we unify all TR tasks and finetune with the pretrained visual encoder.
@@ 3.1 Self-Supervised Pretraining of the Visual Encoder
Image tokens. Define a visual codebook in the latent space $Z\in\mathbb{R}^{K\times D}$, representing $K$ entries of visual tokens $z_{i}\in\mathbb{R}^{D}$. The visual codebook is trained in such a way that an input image, once it is embedded into an image grid, each pixel on the embedded image grid can be substituted with a visual token from the codebook. Decoding this modified image grid will then reconstruct the input image. We use Vector Quantized-Variational AutoEncoder (VQ-VAE[37] to train the visual codebook. Specifically, the tabular image $I$ is tokenized into discrete tokens $z$ after passing through the encoder $q_{\phi}(z|I)$, and the decoder $p_{\psi}(I|z)$ takes these discrete tokens and rebuilds the original image. The training objective of VQ-VAE is to maximize the $\mathbb{E}_{z\sim q_{\phi}(z|I)}[\log p_{\psi}(I|z)]$ with respect to $\phi$ and $\psi$. The training is non-differentialble due to the categorical distribution in selecting visual tokens. Thus, we use Gumbel-Softmax [13] as a reparameterization trick following DALL-E [33]. We have trained the VQ-VAE on 1M and 2M tabular images, where $K=8192$ for 1M, and $K=16384$ for 2M.
Image patches. Given an input tabular image $I\in\mathbb{R}^{H\times W\times C}$, the linear projection divides $I$ into a sequence of flattened 2D patches $I_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $C$ is the number of channels, $(P,P)$ is the size of each image patch, and $N=HW/P^{2}$ is the number of patches. It is implemented by a kernel $P\times P$, stride $P$ convolution. We set $P=16$ and $I\in\mathbb{R}^{448\times 448\times 3}$, thus the sequence length of the image patches is $28\times 28$. Approximately 40% of the sequence is replaced with a masked token and the pretraining objective is to maximize the log-likelihood of the visual tokens of the masked region given the unmasked region. The image tokenizer’s codebook provides the groundtruth visual tokens.
The pretraining task is inspired by the success of masked language [6] and natural image [2] modeling, but our work differs in the following ways:
@@ 3.2 Unified Finetuning Framework
Prior work employed a task-specific decoder for each task, where $\mathbf{S}$ was predicted by an image-to-text pipeline and $\mathbf{B}$ was predicted by a detection head, e.g., Faster R-CNN [36] or DETR [28]. The assumption on predicting $\mathbf{C}$ varies: some assume a PDF is always accompanying the tabular image [28], others rely on external text line detection and text recognition models [40, 12]. We aim to provide a unified task-agnostic training framework, where the input to our model is an image in the form of raw pixels only and the output is text in the form of token sequences. This setting is also generalizable to PDF input modality as we can simply take a screenshot of the PDF. The framework also enables us to leverage the visual encoder pretrained on the unannotated tabular images.
@@
Table cell bbox $\mathbf{B}$. Each cell bbox $\mathbf{b}$ are four continuous coordinates, which are not naturally expressed as discrete tokens. Inspired by Pix2Seq [4], we discretize the coordinate into an integer between 0 and image size. The two directions of an image share the same vocabulary. Since we need to predict all bboxes within a tabular image, each quantized bbox is concatenated together in reading order: from left to right and top to bottom. This formulation completes the quantization and serialization of all bboxes into a sequence of discrete tokens. At inference time, we de-serialize the predicted sequence into groups of four tokens.
Table cell content $\mathbf{C}$. After predicting all bbox coordinates, we only need to perform optical character recognition (OCR) on the image region within each bbox. Note that each cell is defined by a single bbox, thus the cell content can have single or multiple lines of text. In the training stage, the model is trained on a mixture of single line and multi-line dataset. At inference time, we parse all text simultaneously as each cell bbox is independent. Finally, we insert the cell content back into the non-empty cells <td>[]</td> or >[]</td> as the reading order is already preserved in both S and B. We use WordPiece tokenizer [38] with charater-level granularity since OCR requires the model to read instead of understanding the semantics, which significantly reduces the total vocabulary size to less than 6k.
@@
Table 1: UniTable outperforms prior methods and achieves SOTA on four out of five largest publicly available TR datasets across all available tasks. Our method is trained with a task-agnostic language modeling loss and does not rely on external PDF for text extraction and bbox post-processing.